OpenStudy (anonymous):

whats the general answer of y""+2y"+y=0

OpenStudy (anonymous):

homogenous solution \[y^{iv}+2y^{II}+y\] write characteristic equations r^4+2r^2+1=0 solve for r

OpenStudy (anonymous):

how to solve for r in test without calculator

OpenStudy (anonymous):

solve it like you are solving quadratic; try factorng

OpenStudy (anonymous):

i dont think it has any real root.

OpenStudy (anonymous):

(r^2 +1 )(r^2 +1)

OpenStudy (mr.math):

What kind of notation are you using imran, iv and II? :P

OpenStudy (anonymous):

superbowl notation

OpenStudy (anonymous):

y=0 is one solution.. get others by factoring y^3+y+1=0

OpenStudy (anonymous):

it should be sin/cos

OpenStudy (jamesj):

sam, no. And yes, there are no real roots. But that doesn't phase us, because for every complex root, its conjugate is also a root and we know (don't we?) how to put the corresponding solutions back together into real functions.

OpenStudy (anonymous):

yes ..i got u..

OpenStudy (anonymous):

I meant in terms of sin and cos

OpenStudy (jamesj):

the only other thing you need to know is how to deal with repeated roots. For example, what are the solutions of y'' - 2y + y = 0 ? The corresponding characteristic equation is \( r^2 - 2r + 1 =0 \), i.e., r = 1 twice. The general solution is \[ y(x) = c_1e^x + c_2xe^x \]

OpenStudy (anonymous):

ah.... i know that but the real problem in test is finding the roots y=c1Cosx + c2Sinx +x(c3Cosx + c4Sinx)

OpenStudy (anonymous):

that's it , I believe